Crew 315 Operations Report 23-04-2025
SOL: 3
Name of person filing report: Michael Andrews
Non-nominal systems: Charger #7, Space Suit #1, #3, Perseverance Rover, Curiosity Rover
Notes on non-nominal systems:
1. (Unresolved from yesterday) It was noted that Suit #7 wasn’t charging, but further diagnosis from the Commander has isolated the issue to the #7 charger. Troubleshooting this will take place over the next day.
2. Space Suit #3’s neck ring cord broke and needs to be reinstalled. Its cord was cannibalized from Suit #1 to support an immediate EVA and will be fixed this evening. Suit #3’s oxygen cord was also cross-threaded upon install and is stuck. Some quick work with some elbow grease and adjustable pliers will fix it this evening.
3. (Unresolved from yesterday) The Perseverance Rover has two minor issues: the parking brake does not fully engage and have much effect, and the ignition block rotates fully if you turn the key too hard by accident.
4. The Curiosity Rover has a lagging battery indicator. On EVAs, its battery life percentage will drop by only a few points and register much later on the following EVA. This is believed to be fixed by watering the battery with DI water, an action for Mission Support.
ROVERS
Spirit rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: No
Opportunity rover used: Yes
Hours: 325.1 hours (1.3 hours used)
Beginning charge: 100%
Ending charge: 71%
Currently charging: No
Curiosity rover used: Yes
Hours: 313.5 hours (1.7 hours used)
Beginning charge: 100%
Ending charge: 67%
Currently charging: No
Perseverance rover used: Yes
Hours:
Beginning charge:
Ending charge:
Currently charging: No
General notes on rovers: I performed a power study on all three rovers based on 6 EVAs of data. Perseverance has the best battery life, and Curiosity performs equally with Opportunity. However, until Curiosity’s battery is watered, its battery level does not update real-time.
Summary of Hab operations: The crew performed another 100cameras module after breakfast, cooked bread for lunch and burritos for dinner, and performed two EVAs.
Water Use
Time of measurements: 1930
1) Per formula: 41.66 gallons
2) Smart Home Dashboard: 46.5 gallons
Water (static tank, remaining gallons): 390.71 gallons
Static tank pipe heater: Off
Static tank heater: Off
Toilet tank emptied: Yes
Summary of internet: Used to complete reports, learn languages, share team photos via Google drive, researching projects
Summary of suits and radios:
· EVA 5: Suits 8,9 and Radios 3, 4
· EVA 6: Suits 2, 7, 10 and Radios 1, 2, 5
Summary of GreenHab operations: Watered all crops, cultivated fungal project, and opened door when temperatures exceeded 94 F.
WATER USE: 11.0 gallons
Heater: Yes
Supplemental light: 0000 to 0900 and then 1900 to 2200 (light was on all night due to script needing to be rebooted)
Harvest (name, weight in grams): tomatoes (66g), chives (8g), parsley (5g)
Summary of ScienceDome operations: It was a late night last night to map the second rock sample with the 3D scanner, but it proved to be successful and very informative. This morning a third sample was scanned in about an hour… half the time of previous samples! The Crew Journalist also began her desiccated soil experiment in the ScienceDome this afternoon. The fan that cools the power converter was turned on to level 3.
Dual split (Heat or AC, On or Off): Off
Summary of RAM operations:
· Space Suit Repairs: As mentioned above, neck ring cords were swapped on suits 3 and 1 to facilitate EVA 6.
· Crew Artist’s Crazy Straw: Gagnon requested the ability to drink from his water bottle on EVAs without taking off his helmet. I fashioned a clean rubber hose from the science lab into a drinking apparatus that tapes to the suit helmet. With a little bit of adjustment, it worked well on EVA 6. Once the mission is over, it will be cleaned and returned to the dome.
· Tunnel Zip Ties: Only one zip tie was missing after yesterday’s evening gusts! A more thorough check will be performed in the AM. One zip tie was taken for install in next daylight.
· GreenHab Scale: a scale was repurposed from the ScienceDome to measure GreenHab harvests. Still troubleshooting the original scale to see if it can be fixed in the RAM.
· Other Tools Uses: Additional jars (2) were taken for rock specimens. A tool bag, carabiner, and rope tape measurer were taken for the Crew Journalist’s project, and a Rock Hammer was taken for rock specimens. All will be kept in the hab for EVA use until mission duration.
Summary of any observatory issues: None
Summary of health and safety issues: None
Questions, concerns and requests to Mission Support: Great work on fixing the main generator Mission Support! This will lower noise levels and allow us to monitor station health via Home Assistant more consistently.